OpenStudy (anonymous):

Determine whether the point, (3,-2) is a solution to 3x-2y=5 A. Yes B. No

Vocaloid (vocaloid):

let x = 3 and y = -2 then calculate 3x - 2y

OpenStudy (sbuck98):

go to desmos.com

OpenStudy (sbuck98):

it helps

OpenStudy (anonymous):

So yes

OpenStudy (animalover02):

the answer is a

OpenStudy (animalover02):

please fan and metal please

Vocaloid (vocaloid):

3(3)-2(-2) = ?

OpenStudy (animalover02):

its the first answer. look it up

Vocaloid (vocaloid):

nope... 3(3)-2(-2) = 9 + 4 which is not 5.... so the answer is no...

OpenStudy (animalover02):

type the equation in on google. youll get the first as a answer